OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-7, -5) and has slope -1/7. Please show all of your work. Please show all of your work.

OpenStudy (anonymous):

Given: (-7, -5) and m=-1/7 Slope-Intercept Formula: (x-x')=m(y-y') x+7=(-1/7)(y+5) multiply by 7 to get rid of fraction 7x+49=-1(y+5) 7x+49=-y-5 move the y to left side y+7x+49=-5 subt. 49 y+7x=-54 switch x y=-7x-54